OpenStudy (anonymous):

What do you think it means when someone tells you the probability of an event is equal to 0? What is the percentage chance of this event happening?

OpenStudy (anonymous):

If the probability of an event is equal to 0 then it means that given your probability space it will never happen, the percentage would be obviously 0%

OpenStudy (anonymous):

thanks .. can I ask you 2 more

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

We can write a square root as a power of 0.5 (or one half). For example, (32)^0.5 = 5.66. Try this with your calculator! a. Use this notation of a fractional power and a property of exponents (from Unit 1) to prove the following: This is challenging, but try it! √ a * b = √a * b √